A student drops a water balloon out of a dorm window 79 m above the ground. What is the speed in m/s when it hits the ground? Round to the closest whole number.
Answer #1

using 3rd equation of motion

v^2 = u^2 + 2 g h

v^2 = 0 + 2 * 9.8* 79

v = 39 m/s

=========

Comment before rate in case any doubt, will reply for sure.. goodluck